### v1

# Date Fragments: A Hidden Bottleneck of Tokenization for Temporal Reasoning

Abstract
Modern BPE tokenizers often split calendar dates into meaningless fragments, e.g., “20250312” $\rightarrow$ “202”, “503”, “12”, inflating token counts and obscuring the inherent structure needed for robust temporal reasoning. In this work, we (1) introduce a simple yet interpretable metric, termed date fragmentation ratio, that measures how faithfully a tokenizer preserves multi-digit date components; (2) release DateAugBench, a suite of 6500 examples spanning three temporal reasoning tasks: context-based date resolution, format-invariance puzzles, and date arithmetic across historical, contemporary, and future regimes; and (3) through layer-wise probing and causal attention-hop analyses, uncover an emergent date-abstraction mechanism whereby large language models stitch together the fragments of month, day, and year components for temporal reasoning. Our experiments show that excessive fragmentation correlates with accuracy drops of up to 10 points on uncommon dates like historical and futuristic dates. Further, we find that the larger the model, the faster the emergent date abstraction that heals date fragments is accomplished. Lastly, we observe a reasoning path that LLMs follow to assemble date fragments, typically differing from human interpretation (year $\rightarrow$ month $\rightarrow$ day).
Date Fragments: A Hidden Bottleneck of Tokenization
for Temporal Reasoning
Gagan Bhatia1  Maxime Peyrard2  Wei Zhao1 1University of Aberdeen   2Université Grenoble Alpes & CNRS {g.bhatia.24,wei.zhao}@abdn.ac.uk
1 Introduction
Understanding and manipulating dates is a deceptively complex challenge for modern large language models (LLMs). Unlike ordinary words, dates combine numeric and lexical elements in rigidly defined patterns—ranging from compact eight-digit strings such as 20250314 to more verbose forms like “March 14, 2025” or locale-specific variants such as “14/03/2025.” Yet despite their structured nature, these date expressions often fall prey to subword tokenizers that fragment them into semantically meaningless pieces. A tokenizer that splits “2025-03-14” into “20”, “25”, “-0”, “3”, “-1”, “4” not only inflates the token count but also severs the natural boundaries of year, month, and day. This fragmentation obscures temporal cues and introduces a hidden bottleneck: even state-of-the-art LLMs struggle to resolve, compare, or compute dates accurately when their internal representations have been so badly fragmented. This issue is critical for real-world applications:
Figure 1: Internal processing of dates for temporal reasoning. Here F=0.4 shows the date fragmentation ratio.
Mis-tokenized dates can undermine scheduling and planning workflows, leading to erroneous calendar invites or appointments Vasileiou and Yeoh (2024). They can skew forecasting models in domains ranging from time-series analysis Tan et al. (2024); Chang et al. (2023) to temporal knowledge graph reasoning Wang et al. (2024). In digital humanities and historical scholarship, incorrect splitting of date expressions may corrupt timelines and misguide interpretative analyses Zeng (2024). As LLMs are increasingly deployed in cross-temporal applications, such as climate projectionWang and Karimi (2024), economic forecasting Carriero et al. (2024); Bhatia et al. (2024), and automated curriculum scheduling Vasileiou and Yeoh (2024), the brittleness introduced by subword fragmentation poses a risk of propagating temporal biases and inaccuracies into downstream scientific discoveries and decision-making systems Tan et al. (2024).
In this work, we provide a pioneer outlook on the impact of date tokenization on downstream temporal reasoning. Figure 1 illustrates how dates are processed internally for temporal reasoning.
Our contributions are summarized as follows:
  1. (i)
    We introduce DateAugBench, a benchmark dataset comprising 6,500 examples with 21 date formats. It is leveraged to evaluate a diverse array of LLMs from 8 model families in three temporal reasoning tasks.
  2. (ii)
    We present date fragmentation ratio, a metric that measures how fragmented the tokenization outcome is compared to the actual year, month, and day components. We find that the fragmentation ratio generally correlates with temporal reasoning performance, namely that the more fragmented the tokenization, the worse the reasoning performance.
  3. (iii)
    We analyse internal representations by tracing how LLMs “heal” fragmented date embeddings in their layer stack—an emergent ability that we term date abstraction. We find that larger models quickly can compensate for date fragmentation to achieve high accuracy of temporal reasoning at early layers.
  4. (iv)
    We leverage causal analysis to interpret how LLMs stitch date fragments for temporal reasoning. Our results show that LLMs follow a reasoning path that is typically not aligned with human interpretation (year $\rightarrow$ month $\rightarrow$ day), but relies on subword fragments that statistically represent year, month, and day, and stitch them in a flexible order that is subject to date formats.
Our work fills the gap between tokenisation research Goldman et al. (2024); Schmidt et al. (2024) and temporal reasoning Su et al. (2024); Fatemi et al. (2024), and we suggest future work to consider date-aware vocabularies and adaptive tokenizers to ensure that date components remain intact.
2 Related Works
Tokenisation as an information bottleneck.
[…]
Temporal reasoning in large language models.
Despite rapid progress on chain-of-thought and process-supervised reasoning, temporal cognition remains a conspicuous weakness of current LLMs. Benchmarks such as TimeBench Chu et al. (2024), TempReason Tan et al. (2023), Test-of-Time Fatemi et al. (2024), MenatQA Wei et al. (2023) and TimeQA Chen et al. (2021) reveal large gaps between model and human performance across ordering, arithmetic and co-temporal inference. Recent modelling efforts attack the problem from multiple angles: temporal-graph abstractions Xiong et al. (2024), instruction-tuned specialists such as Timo Su et al. (2024), pseudo-instruction augmentation for multi-hop QA Tan et al. (2023), and alignment techniques that re-ground pretrained models to specific calendar years Zhao et al. (2024). Yet these approaches assume a faithful internal representation of the input dates themselves. By introducing the notion of date fragmentation and demonstrating that heavier fragmentation predicts up to ten-point accuracy drops on DateAugBench, we uncover a failure mode that is orthogonal to reasoning algorithms or supervision: errors arise before the first transformer layer, at the level of subword segmentation. Addressing this front-end bottleneck complements, rather than competes with, existing efforts to improve temporal reasoning in LLMs.
3 DateAugBench
Dataset and Task # Formats # Raw Size Evaluation
Example GT
Context based 6 500 3000 Which team did Omid Namazi play for in 06/10/1990? Maryland Bays
Date Format Switching 10 150 1500 Are 20251403 and March 14th 2025 referring to the same date? Yes
Date Arithmetic 5 400 2000 What date is 10,000 days before 5/4/2025? 18 November 1997; 17 December 1997
Total 21 1500 6500
Table 1: Overview and examples of task splits in DateAugBench.
We introduce DateAugBench, benchmark designed to isolate the impact of date tokenisation on temporal reasoning in LLMs. DateAugBench comprises 6,500 augmented examples drawn from two established sources, TimeQA Chen et al. (2021) and TimeBench Chu et al. (2024), distributed across three tasks splits (see Table 1). Across all the splits, our chosen date formats cover a spectrum of common regional conventions (numeric with slashes, dashes, or dots; concatenated strings; two-digit versus four-digit years) and deliberately introduce fragmentation for atypical historical (e.g. “1799”) and future (e.g. “2121”) dates. This design enables controlled measurement of how tokenization compression ratios and subsequent embedding recovery influence temporal reasoning performance.
Context-based task.
In the Context-based split, we sample 500 question–context pairs from TimeQA, each requiring resolution of a date mentioned in the passage (e.g. Which team did Omid Namazi play for in 06/10/1990?). Every date expression is systematically rendered in six canonical serialisations—including variants such as MM/DD/YYYY, DD-MM-YYYY, YYYY.MM.DD and concatenations without delimiters—yielding 3,000 examples that jointly probe tokenisation fragmentation and contextual grounding.
Simple Format Switching task.
The Simple Format Switching set comprises 150 unique date pairs drawn from TimeBench, posed as binary same-day recognition questions (e.g. “Are 20251403 and 14th March 2025 referring to the same date?”). Each pair is presented in ten different representations, spanning slash-, dash-, and dot-delimited formats, both zero-padded and minimally notated, to stress-test format invariance under maximal tokenisation drift. This produces 1,500 targeted examples of pure format robustness. We also have examples where the dates are not equivalent, complicating the task.
Date Arithmetic task.
The Date Arithmetic split uses 400 arithmetic instances from TimeBench (e.g. What date is 10,000 days before 5/4/2025?). With the base date serialised in five distinct ways—from month-day-year and year-month-day with various delimiters to compact eight-digit forms. This results in 2,000 examples that examine the model’s ability to perform addition and subtraction of days, weeks, and months under various token fragmentation.
4 Experiment Design
4.1 Date Tokenization
Tokenizers.
For tokenization analysis, we compare a deterministic, rule-based baseline tokenizer against model-specific tokenizers. The baseline splits each date into its semantic components—year, month, day or Julian day—while preserving original delimiters. For neural models, we invoke either the OpenAI tiktoken encodings (for gpt-4, gpt-3.5-turbo, gpt-4o, text-davinci-003) or Hugging Face tokenizers for open-source checkpoints. Every date string is processed to record the resulting sub-tokens, token count, and reconstructed substrings.
Distance metric.
To capture divergence from the ideal, we define a distance metric $\theta$ between a model’s token distribution and the baseline’s:
$\theta(\mathbf{t},\mathbf{b})=1-\frac{\mathbf{t}\cdot\mathbf{b}}{|\mathbf{t}|,|\mathbf{b}|},$ (1)
where $\mathbf{t}$ and $\mathbf{b}$ are vectors of sub-token counts for the model and baseline, respectively. A larger $\theta$ indicates greater sub-token divergence.
Date fragmentation ratio.
Building on $\theta$, we introduce the date fragmentation ratio $F$, which quantifies how fragmented a tokenizer’s output is relative to the baseline. We initialise $F=0.0$ for a perfectly aligned segmentation and apply downward adjustments according to observed discrepancies: a 0.10 penalty if the actual year/month/day components are fragmented (i.e., $\mathbf{1}_{\mathrm{split}}=1$) , a 0.10 penalty if original delimiters are lost (i.e., $\mathbf{1}_{\mathrm{delimiter}}=1$), a 0.05 penalty multiplied by the token count difference $(N-N_{b}\bigr{)}$ between a tokenizer and the baseline, and a $0.30\times\theta$ penalty for distributional divergence. The resulting $F\in[0,1]$ provides an interpretable score: values close to 0 denote minimal fragmentation, and values near 1 indicate severe fragmentation.
$F$ $=0.10\times\mathbf{1}_{\mathrm{split}}+0.10\times\mathbf{1}_{\mathrm{delimiter}}$ (2)
$\quad+0.05\times\bigl{(}N-N_{b}\bigr{)}+0.30\times\theta$
This date fragmentation ratio is pivotal because tokenisation inconsistencies directly impair a model’s ability to represent and reason over temporal inputs. When date strings are split non-intuitively, models face inflated token sequences and fragmented semantic cues, potentially leading to errors in tasks such as chronological comparison, date arithmetic, and context-based resolution. By quantifying fragmentation explicitly through $F$, we reveal hidden limitations in existing tokenizers, inform selections of robust architectures for time-sensitive applications.
4.2 Temporal Reasoning Evaluation
Models.
We evaluate a spectrum of model ranging from 0.5 B to 14 B parameters: five open-source Qwen 2.5 models (0.5 B, 1.5 B, 3 B, 7 B, 14 B) Yang et al. (2024), two Llama 3 models (3 B, 8 B) Touvron et al. (2023b), and two OLMo Groeneveld et al. (2024) models (1 B, 7 B). For comparison with state-of-the-art closed models, we also query the proprietary GPT-4o and GPT-4o-mini endpoints via the OpenAI API OpenAI et al. (2024).
LLM-as-a-judge.
To measure how date tokenization affects downstream reasoning, we employ an LLM-as-judge framework using GPT-4o. For each test instance in DateFragBench, we construct a JSONL record that includes the question text, the model’s predicted answer, and a set of acceptable gold targets to capture all semantically equivalent date variants (e.g., both “03/04/2025” and “April 3, 2025” can appear in the gold label set). This record is submitted to GPT-4o via the OpenAI API with a system prompt instructing it to classify the prediction as CORRECT, INCORRECT, or NOT ATTEMPTED. A prediction is deemed CORRECT if it fully contains any one of the gold target variants without contradiction; INCORRECT if it contains factual errors relative to all gold variants; and NOT ATTEMPTED if it omits the required information. We validate GPT-4o’s reliability by randomly sampling 50 judged instances across all splits and obtaining independent annotations from four human reviewers. GPT-4o’s classifications agree with the human consensus on 97% of cases, yielding a Cohen’s $\kappa$ of 0.89, which affirms the reliability of our automated evaluation.
4.3 Internal Representations
Figure 2: Illustration of how LLMs with various model sizes process dates. TCP means Tokenization Compensation Point, defined as the first layer at which LLMs achieve above-chance accuracy (see details in Sec. 6).
Layerwise probing.
We use four Qwen2.5 Yang et al. (2024) model checkpoints (0.5B, 1.5B, 3B, and 7B parameters) to trace how temporal information is processed internally across different layers. During inference, each question is prefixed with a fixed system prompt and a chain-of-thought cue, then passed through the model in evaluation mode. At each layer $i$, we extract the hidden-state vector corresponding to the final token position, yielding an embedding $h_{i}\in\mathbb{R}^{d}$ for that layer. Repeating over all examples produces a collection of layer-wise representations for positive and negative cases. We then quantify the emergence of temporal reasoning by training lightweight linear probes on these embeddings. For layer $i$, the probe is trained to distinguish “same-date” vs “different-date” examples. To explain when the model’s date understanding is achieved, we define the tokenization compensation point as the layer at which the model’s representation correctly represents the date in the given prompt. We experiment with this idea across various model sizes, aiming to test our hypothesis: larger models would recover calendar-level semantics from fragmented tokens at earlier stages, i.e., tokenization compensation is accomplished at early layers, as illustrated in Figure 2.
Causal attention-hop analysis.
We introduce a framework intended to understand in which order date fragments are stitched together for LLMs to answer a temporal question. Figure 1 depicts the idea of our framework: given an input prompt requiring a date resolution (e.g., “Is 28052025 the same date as 28th of May 2025?”), we define two sets of tokens: (1) concept tokens corresponding to year, month, and day fragments, and (2) decision tokens corresponding to the model answer (“yes” or “no”). Our framework aims to identify a stitching path for temporal reasoning, or reasoning path for short. A reasoning path is defined as a sequence of tokens containing date fragments and the model answer [^1] The idea of reasoning paths was introduced by Lindsey et al. (2025), which we leverage to interpret how LLMs address date fragments for temporal reasoning. . Given that there are multiple potential paths, we score each path and select the highest-scoring one as the LLM’s reasoning path for the given prompt. To score a reasoning path, our idea is the following: we identify when a date fragment or model answer is activated, by which input token and at which layer, and then determine how important each input token is for the date fragment and model answer. Our idea is implemented by using two different approaches: (i) next token prediction (§A.2.1): how likely a date fragment and model answer follows a given input token and (ii) token importance (§A.2.2): how important an input token is to a date fragment and model answer by replacing the input token with a random token. Lastly, we combine the results of the two approaches to yield the final score of a reasoning path (§A.2.3). This causal framework not only pinpoints where and when date fragments are activated, but also in which order they are stitched together to yield the model answer.
5 Experiment Results
5.1 Date fragmentation
Model Past Near Past Present Future Avg
Baseline 0.00 0.00 0.00 0.00 0.00
OLMo 0.15 0.14 0.07 0.25 0.15
GPT-3 0.17 0.14 0.06 0.25 0.16
Llama 3 0.29 0.28 0.27 0.30 0.29
GPT-4o 0.32 0.31 0.22 0.30 0.29
GPT-3.5 0.47 0.22 0.26 0.36 0.33
GPT-4 0.36 0.26 0.29 0.39 0.33
Qwen 0.58 0.55 0.49 0.58 0.55
Gemma 0.58 0.55 0.49 0.58 0.55
DeepSeek 0.58 0.55 0.49 0.58 0.55
LlaMa 0.63 0.63 0.63 0.63 0.63
Phi 0.63 0.63 0.63 0.63 0.63
Table 2: Date fragmentation ratio across models and data splits over time. Here we refer to the family of models that use the same tokenizers.
Models Context Rlt Fmt Switch Date Arth. Avg.
GPT-4o-mini 53.20 95.66 56.67 68.51
OLMo-2-7B 32.13 97.24 64.72 64.70
Qwen2.5 14B 47.56 94.56 51.35 64.49
Qwen2.5 7B 39.56 91.24 40.56 57.12
Qwen2.5 3B 25.45 90.10 39.45 51.67
LLama3.1 8B 26.20 90.22 34.50 50.31
Qwen2.5 1.5B 21.32 89.65 32.34 47.77
Qwen2.5 0.5B 10.23 88.95 31.32 43.50
OLMo-2-1B 9.26 90.09 25.90 41.75
LLama3.2 3B 9.51 88.45 23.66 40.54
Table 3: Average accuracies per task. Context Rlt stands for context based resolution task, Fmt Switch refers to the format switching task, and Date Arth. refers to the date arithmetic task.
Cross-temporal performance.
Table 2 reports the mean date fragmentation ratio across four temporal regimes—Past (pre–2000), Near Past (2000–2009), Present (2010–2025), and Future (post–2025)—for each evaluated model. A ratio of 0.00 signifies perfect alignment with our rule-based baseline tokenizer, whereas higher values indicate progressively greater fragmentation. The rule-based Baseline unsurprisingly attains the maximal ratio of 0.00 in all periods, serving as a lower bound. Among neural architectures, OLMo Groeneveld et al. (2024) demonstrates the highest robustness, with an average fragmentation ratio of 0.15, closely followed by GPT-3 at 0.16. Both maintain strong fidelity across temporal splits, although performance dips modestly in the Future category (0.25), reflecting novel token sequences not seen during pre-training.
Model Tokenized output Frag-ratio
Baseline 10 27 1606 0.00
OLMo 10 27 16 06 0.34
Llama 3 102 716 06 0.40
GPT-3 1027 16 06 0.40
GPT-4o 102 716 06 0.40
Gemma 1 0 2 7 1 6 0 6 0.55
DeepSeek 1 0 2 7 1 6 0 6 0.55
Cohere 1 0 2 7 1 6 0 6 0.55
Qwen 1 0 2 7 1 6 0 6 0.55
Phi 3.5 _ 1 0 2 7 1 6 0 6 0.60
Llama 2 _ 1 0 2 7 1 6 0 6 0.60
Table 4: Tokenisation of the MMDDYYYY string ‘‘10271606’’ across models.
Impact of subtoken granularity.
A closer look, from Table 4, at sub-token granularity further explains these trends. Llama 3 Touvron et al. (2023b) and the GPT OpenAI et al. (2023) families typically segment each date component into three-digit sub-tokens (e.g., “202”, “504”, “03”), thus preserving the semantic unit of “MMDDYYYY” as compact pieces. OLMo Groeneveld et al. (2024) splits the date tokens into two digit tokens (e.g., “20”, “25”). By contrast, Qwen Yang et al. (2024) and Gemma Team et al. (2024) models break dates into single-digit tokens (e.g., “2”, “5”), whereas Phi Abdin et al. (2024) and LLama Touvron et al. (2023a) divide it into single-digit tokens with an initial token (e.g. “_”, “2”, “0”, “2”, “5”), inflating the token count. Although single-digit tokenisation can enhance models’ ability to perform arbitrary numeric manipulations (by treating each digit as an independent unit), it comes at the expense of temporal abstraction: the tight coupling between day, month, and year is lost, inflating the compression penalty and increasing the $\theta$ divergence from the baseline.
5.2 DateFragBench Evaluation
Figure 3: Date fragmentation ratio versus date resolution accuracy, stratified by temporal regime and six LLMs: OLMo, Llama 3, GPT-4o, Qwen, Gemma, Phi.
Figure 4: Date fragmentation ratio versus date resolution accuracy, stratified by six formats and six LLMs.
Performance on temporal reasoning tasks.
We compare model accuracies in three tasks: Context-based Resolution, Format Switching, and Date Arithmetic (see the results in Table 3). All models effectively solve Format Switching (e.g. 97.2% for OLMo-2-7B, 95.7% for GPT-4o-mini, 94.6% for Qwen2.5-14B, 90.2% for Llama3.1-8B). By contrast, Context Resolution and Arithmetic remain challenging: GPT-4o-mini scores 53.2% and 56.7%, Qwen2.5-14B 47.6% and 51.4%, Llama3.1-8B 26.2% and 34.5%, and OLMo-2-7B 32.1% and 64.7%, respectively. The fact that arithmetic performance consistently exceeds resolution suggests that, given a correctly tokenized date, performing addition or subtraction is somewhat easier than resolving the date within free text—which requires encyclopedic knowledge.
Correlating date fragmentation with model accuracy over time.
Figure 3 plots date fragmentation ratio against resolution accuracy, with 24 data points across six models and four temporal splits. Accuracy rises as we move from Past (1600-2000) to the Near Past (2000–2009) and peaks in the Present (2010–2025), mirroring the negative correlation between fragmentation and accuracy (dashed line, Pearson correlation of $-0.61$). We note that the correlation is not particularly strong. This is because (i) for some models, their date fragmentation ratios remain unchanged across temporal data splits and (ii) models differ greatly by their sizes: a larger model often outperforms a substantially smaller model in resolution accuracy, even if the former has a much higher fragmentation ratio.
As seen from Table 5, GPT-4o-mini climbs from 61.7 % in the Past to 67.9 % in the Near Past, peaks at 70.5 % for Present, and falls to 58.2 % on Future dates. Qwen-2.5-14B and Llama-3.1-8B trace the same contour at lower absolute levels. OLMo-2-7B shows the steepest Near-Past jump (49.5 $\rightarrow$ 62.4 %) and achieves the highest Present accuracy (73.6 %), consistent with its finer-grained tokenisation of “20XX” patterns. These results indicate that while finer date tokenisation (i.e., lower fragmentation ratios) boosts performance up to contemporary references, today’s models still generalise poorly to genuinely novel (post-2025) dates, highlighting an open challenge for robust temporal reasoning.
Correlating date fragmentation with model accuracy over formats.
Figure 4 plots model accuracy against date fragmentation ratio across six date formats and six LLMs. A moderate negative trend emerges (dashed line, Pearson correlation of $-0.42$): formats that contain explicit separators (DD-MM-YYYY, DD/MM/YYYY, YYYY/MM/DD) are tokenised into more pieces and, in turn, resolved more accurately than compact, separator-free strings (DDMMYYYY, MMDDYYYY, YYYYMMDD). As shown in Table 6, GPT-4o-mini tops every format and receives a moderate performance drop from 71.2 % on DD/MM/YYYY to 61.2 % on DDMMYYYY, with the highest overall average (66.3 %). OLMo-2-7B and Qwen-2.5-14B both exceed 70 % on the highly fragmented YYYY/MM/DD form, but slip into the low 50s on MMDDYYYY and YYYYMMDD. Lower date fragmentation ratio models, such as Llama-3.1-8B and Phi-3.5, lag behind; their accuracy plunges below 40 %. Even so, all models score much better on separator-rich formats compared to the date formats without separators. In summary, model accuracy is correlated to how cleanly a model can tokenize the string into interpretable tokens: more visual structure (slashes or dashes) means lower fragmentation, which suggests more straightforward reasoning, and in turn, leads to better performance.
6 In which layer do LLMs compensate for date fragmentation?
(a) Past
(b) Present
Figure 5: Layer-wise accuracies in the two periods: Past and Present.
Figure 6: Reasoning path for the “03122025 is a valid date” prompt.
Layerwise linear probing.
To pinpoint in which layer a model learns to recognize two equivalent dates, we define the tokenization compensation point (TCP) as the earliest layer at which a lightweight linear probe on the hidden state achieves above-chance accuracy, which is defined as 80%, on the date equivalence task. Figure 5(a) reports TCPs for the Dates_Past benchmark (1600–2010): Qwen2.5-0.5B reaches TCP at layer 12 (50% depth), Qwen2.5-1.5B at layer 15 (53.6%), Qwen2.5-3B at layer 8 (22.2%), and Qwen2.5-7B at layer 4 (14.3%). The leftward shift of the 3B and 7B curves suggests how larger models recover calendar-level semantics from fragmented tokens more rapidly. Figure 5(b) shows the Dates_Present benchmark (2010–2025), where only the 1.5B, 3B, and 7B models surpass TCP—at layers 16 (57.1%), 21 (58.3%), and 17 (60.7%), respectively—while the 0.5B model never does. The deeper TCPs here reflect extra layers needed to recombine the two-digit “20” prefix, which is fragmented unevenly by the tokenizer. In Figure 10, we evaluate Dates_Future (2025–2599), where novel four-digit sequences exacerbate fragmentation. Remarkably, TCPs mirror the Past regime: layers 12, 15, 8, and 4 for the 0.5B, 1.5B, 3B, and 7B models, respectively. This parallelism indicates that model scale dictates how quickly LLMs can compensate for date fragmentation to achieve high accuracy, even when dates are novel.
Tokenization compensation point.
Overall, we observe a sharp decline in TCP as model size increases: small models defer date reconstruction to middle layers, whereas the largest model does so within the first quarter of layers. Across all the three temporal benchmarks, TCP shifts steadily toward the first layers as model size grows.
7 How do LLMs stitch date fragments for temporal reasoning?
Causal path tracing.
To investigate how LLMs like Llama 3 Touvron et al. (2023b) internally stitch date fragments to yield a model answer, we apply our casual framework to identify the model’s reasoning path over a specific prompt. Figure 6 plots model layers on the $y$ axis against prompt tokens (e.g., Is 03122025 a valid date?) on the $x$ axis. Green arrows mark the reasoning path with the highest score that is responsible for generating the answer “yes”. Date fragments “25”, “220”, “031”, and the model answer “yes” are activated in sequence at layer 26-27 by the input tokens “is”, “031”, “a” and “Answer” respectively. As such, the model performs a kind of discrete, step-by-step token aggregation, stitching together substrings of the input until a binary valid/invalid verdict emerges.
Misalignment between LLMs and human.
In contrast, human readers parse dates by immediately mapping each component to a coherent temporal schema: “03” is March, “12” is day of month, “2025” is year, and then checking whether the day falls within the calendar bounds of that month. Humans bring rich world knowledge of calendars and leap-year rules to bear in parallel. However, LLMs exhibit no explicit calendar “module”; instead, they rely on learned statistical associations between digit-patterns and the training-time supervisory signal for “valid date”. The reasoning path in Figure 6 thus illustrates a fundamentally different mechanism of date comprehension in LLMs, based on date fragments re-routing rather than holistic semantic interpretation. We repeated causal tracing on 100 date strings in 6 different date formats to test whether the reasoning path difference between human and LLMs is consistent across date formats. In most of cases, we observe that model reasoning paths are not aligned with human interpretation (year $\rightarrow$ month $\rightarrow$ day), rather rely on sub-word fragments that statistically represent year, month, and day, and stitch these date fragments in a flexible order that is subject to date formats (see examples in Figures 7-8). However, such a reasoning path becomes tricky when a date is greatly fragmented: given the date abstraction is learned from frequency rather than hard-coded rules, the abstraction is biased toward standard Western formats and contemporary years. As a result, a model often addresses popular dates (in the same format) with similar reasoning paths. However, the reasoning path becomes obscure on rare, historical, or locale-specific strings outside the distribution of pre-training data (see Figure 9).
8 Conclusion
In this paper, we identified date tokenization as a critical yet overlooked bottleneck in temporal reasoning with LLMs. We demonstrated a correlation between date fragmentation and task performance in temporal reasoning, i.e., the more fragmented the tokenization, the worse the reasoning performance. Our layerwise and causal analyses in LLMs further revealed an emergent “date abstraction” mechanism that explains when and how LLMs understand and interpret dates. Our results showed that larger models can compensate for date fragmentation by stitching fragments into a unified “date” concept, while the stitching process appears to be accomplished via a reasoning path that connects date fragments in a flexible order, differing from human interpretation from year to month to day.
Limitations
While our work demonstrates the impact of date tokenization on LLMs for temporal reasoning, there are several limitations. First, DateAugBench focuses on a finite set of canonical date serialisations and does not capture the full diversity of natural-language expressions (e.g., “the first Monday of May 2025”) or noisy real-world inputs like OCR outputs. Second, our experiments evaluate a representative but limited pool of tokenizers and model checkpoints (up to 14B parameters); therefore, the generalizability of date fragmentation ratio and our probing and causal analyses to very large models with 15B+ parameters remains unknown. Third, while the fragmentation ratio measures front-end segmentation fidelity, it does not account for deeper world-knowledge factors such as leap-year rules, timezone conversions, and culturally grounded calendar systems, all of which would influence temporal interpretation. Fourth, though we introduce a date fragmentation ratio metric to characterise how tokenisation affects reasoning, we have not yet evaluated this metric.
Lastly, the core idea of our causal framework is inspired by Lindsey et al. (2025); however, our extension to temporal reasoning is not evaluated. Future work should extend to more diverse date expressions, broader model and tokeniser families, equipping tokenisers with external calendar-wise knowledge to improve further robust temporal reasoning, and conducting rigorous evaluation of the fragmentation ratio metric and the causal framework.
Ethical Considerations
DateAugBench is derived solely from the public, research-licensed TimeQA and TimeBench corpora that do not contain sensitive data; our augmentation pipeline rewrites only date strings. However, our dataset focuses on 21 Anglo-centric Gregorian formats. Therefore, our data potentially reinforce a Western default and overlook calendars or numeral systems used in many other cultures, and our date fragmentation metric may over-penalise tokenisers optimised for non-Latin digits.
Acknowledgements
We gratefully thank Madiha Kazi, Cristina Mahanta, and MingZe Tang for their support of conducting human evaluation for LLMs-as-judge.
Appendix A Appendix
A.1 Experiment Design
Implementation details of evaluation.
The evaluation pipeline is implemented in Python and supports asynchronous API requests with retry logic, as well as multiprocessing to handle thousands of examples efficiently. After collecting GPT-4o’s label for each instance, we map CORRECT/INCORRECT NOT ATTEMPTED to categorical scores A, B, and C. We then compute three core metrics: overall accuracy (proportion of A scores), given-attempted accuracy (A over A+B), and the F1 score, defined as the harmonic mean of overall and given-attempted accuracy. Results are reported both globally and stratified by task split (Context-based, Format Switching, Date Arithmetic) and by temporal category (Past, Near Past, Present, Future). We adopt the sample prompts introduced in SimpleQA Wei et al. (2024) as our LLM-as-judge queries, ensuring consistent scoring instructions across all evaluations. Our specific prompt used for evaluation can be found in Table 7. We have presented our examples of LLM as a judge and human evaluation in Table 8.
Date ambiguities.
We explicitly enumerate all valid variants in the gold label set for each example to handle multiple correct answers arising from date-format ambiguities. This ensures that any prediction matching one of these variants is marked correct, avoiding penalisation for format differences.
Synthetic benchmark construction for linear probing.
We construct a suite of synthetic true–false benchmarks to isolate temporal reasoning across different reference frames. For the Dates_Past, Dates_Present, and Dates_Future datasets, we sample 1,000 date–date pairs each, drawing calendar dates uniformly from the appropriate range and rendering them in two randomly chosen, distinct formatting patterns (Ymd vs d/m/Y). Exactly half of each set are “YES” examples (identical dates under different formats), which are our positive examples, and half are “NO” (different dates), which are our negative examples. All three datasets are balanced, shuffled, and split into equal positive and negative subsets to ensure fair probing.
A.2 Causal Attention–Hop Analysis
A.2.1 Next Token Prediction
We treat each token in the prompt as a candidate “concept” to follow. After the model processes the input, it produces a hidden vector $h_{\ell,p}$ per token at position $p$ and layer $\ell$ To see how likely a concept $c$ (e.g., a date fragment and model answer) follows each input token, we project $h_{\ell,p}$ through $W_{U}$ to yield the “probability” distribution of vocabulary tokens, and denote $s^{c}_{\ell,p}$ as the “probability” of the concept being the next token.
$z_{\ell,p}\;=\;W_{U}\,h_{\ell,p},\qquad s^{c}_{\ell,p}\;=\;z_{\ell,p}[t_{c}],$ (1)
where $t_{c}$ is the index of concept $c$ in the vocabulary.
A.2.2 Token Importance
To measure how important an input token is to a concept (e.g., a date fragment and model answer), we replace the token with an unrelated one (e.g., “Dallas” $\rightarrow$ “Chicago”) and compute the probability drop of the concept incurred by the replacement, denoted as $I_{c,p}$ (which we compute only at the last layer):
$I_{c,p}\;=\;\sigma\bigl{(}z_{p}[t_{c}]\bigr{)}\;-\;\sigma\bigl{(}\tilde{z}_{p}[t_{c}]\bigr{)},$ (2)
where $\sigma$ is a softmax function. The bigger the $I_{c,p}$, the more important the original token at position $p$ for the concept $c$.
A.2.3 Path scoring
A reasoning path $\mathcal{P}=(c_{1},\dots,c_{k})$ is a sequence of tokens, indicating in which order date fragments are stitched together for LLMs to answer a temporal question. We score each potential path by blending five components (ordering, activation strength, causal strength, gap penalty, and confidence in the final concept), into a single score:
$S(\mathcal{P})$ $=\;\alpha\times S_{\text{order}}+\beta\times S_{\text{act}}+\gamma\times S_{\text{causal}}$
$\quad-\eta\times S_{\text{gap}}+\kappa\times S_{\text{final}}$ (1)
Each term is designed to reward a different desirable property:
  • Ordering: we give points if the concepts appear in roughly left-to-right order in the prompt, and secondarily in increasing layer order:
    $S_{\text{order}}=0.7\times\mathbf{1}[p_{1}\leq\cdots\leq p_{k}]$
    $+0.3\times\mathbf{1}[\ell_{1}\leq\cdots\leq\ell_{k}],$ (2)
    where $\mathbf{1}$ is an indicator function, $p_{i}=\max_{\ell,p}s^{c_{i}}_{\ell,p}$ indicating the position of the most important input token for a concept $c_{i}$ at the last layer. Similarly, $\ell_{i}$ is the layer at which an input token pays the most attention to the concept $c_{i}$.
  • Activation: we compute the average position of the most important input token for a concept from 1 to $k$, and normalize by a threshold $\tau=0.2$, and clip to 1:
    $S_{\text{act}}=\min\!\bigl{(}\tfrac{1}{k}\sum_{i=1}^{k}p_{i}\,/\,\tau,\;1\bigr{)},$
  • Causal strength: we use the token importance score, denoted as $d_{i}=|I_{c_{i+1},p_{i}}|$ between two adjacent concepts $c_{i+1}$ and $c_{i}$, upweight latter scores, and downweight missing links by a coverage term $\rho$, which is defined as the fraction of actual causal connections observed between consecutive concepts out of the total possible consecutive pairs in the path. The combined score then multiplies the weighted average of the $d_{i}$ by $\tfrac{1}{2}+\tfrac{1}{2}\rho$, giving:
    $S_{\text{causal}}=\bigl{(}\tfrac{\sum_{i}w_{i}d_{i}}{\sum_{i}w_{i}}\bigr{)}\bigl{(}0.5+0.5\rho\bigr{)},$
    where $w_{i}=0.5+0.5\frac{i-1}{k-2}$.
  • Gap penalty: to discourage large jumps in position, we compute the mean gap $\bar{g}$ and apply a small multiplier $\lambda=0.1$:
    $S_{\text{gap}}=1-\lambda\,\bar{g},\quad S_{\text{gap}}\leq 1.$
    This is done to encourage model paths to think step by step instead of directly jumping to the conclusion (yes/no)
  • Final confidence: We compute the position of the most important input token for the last concept $c_{k}$:
    $S_{\text{final}}=\max_{\ell,p}s^{c_{k}}_{\ell,p}.$
The reasoning path with the highest total score $S(\mathcal{P})$ is chosen as the model’s reasoning path over a specific prompt. We note that Ordering, Activation, Gap penalty and Final confidence components are built upon next token prediction signals $s^{c}_{\ell,p}$, whereas the Causal strength component is derived solely from token importance score $I_{c_{i+1},\pi_{i}},$ i.e. the drop in the softmax probability for concept $c_{i+1}$ when the token at position $p_{i}$ is replaced.
Models Past Near Past Present Future
GPT-4o-mini 61.66 67.93 70.51 58.23
OLMo-2-7B 49.45 62.35 73.56 43.45
Qwen2.5 14B 58.97 64.80 67.22 55.69
Qwen2.5 7B 51.41 55.98 57.98 48.55
Qwen2.5 3B 46.50 50.25 51.98 43.91
LLama3.1 8B 45.28 48.82 50.48 42.76
Qwen2.5 1.5B 42.99 46.16 47.69 40.60
Qwen2.5 0.5B 39.15 41.68 43.00 36.98
OLMo-2-1B 36.07 38.09 40.49 34.07
LLama3.2 3B 36.48 38.57 39.74 34.46
Table 5: Model accuracy on context-based questions across four data splits over time.
Model DD-MM-YYYY DD/MM/YYYY YYYY/MM/DD DDMMYYYY MMDDYYYY YYYYMMDD Avg.
OLMo 64.70 64.56 65.35 52.35 54.56 50.41 58.65
Llama 3 50.31 50.89 53.45 38.45 40.24 34.56 44.65
GPT-4o 68.51 71.23 69.24 61.23 62.34 64.98 66.25
Qwen 64.49 62.35 73.56 46.50 50.25 51.98 58.19
Gemma 58.90 58.97 64.80 47.22 46.50 50.25 54.44
Phi 47.23 46.07 48.09 39.15 41.68 43.00 44.20
Table 6: Model accuracy on context-based questions across date formats.
Figure 7: Reasoning path for the “03/12/2025 is a valid date” prompt.
Figure 8: Reasoning path of the “03-12-2025 is a valid date” prompt.
Figure 9: Reasoning path of the “03121325 is a valid date” prompt, where year $=1325$.
Figure 10: Layer-wise accuracies in the Future period
LLM-as-Judge Evaluation Prompt Your task: Evaluate one prediction at a time. You receive: Question – the task prompt shown to the model Gold targetall answers that are considered correct Predicted answer – the model’s response Return one letter only: A CORRECT prediction fully matches one gold variant B INCORRECT prediction contradicts or misses required info C NOT_ATTEMPTED prediction refuses, guesses, or answers irrelevantly General rules: 1. Match semantics, ignore capitalisation, punctuation, order. 2. If any statement contradicts the gold target, grade B. 3. Hedging ("I think…") is fine if the correct info is present and no incorrect info is added. 4. Partial answers are B. Typos that preserve meaning are allowed. DateAugBench specifics: Date format ambiguity: gold lists every valid interpretation; accept any. Date arithmetic: prediction must match day, month, year of a listed variant, any textual format allowed. Format-switch questions: answer with any synonym of Yes/True or No/False. Numeric answers – must match the gold number to the last shown significant digit. Output format
Return exactly one capital letter:
A   or   B   or   C No additional text or punctuation. Example template
Question: {question}
Gold target: {target}
Predicted answer: {predicted_answer}
Now grade: A   or   B   or   C
Table 7: LLM-as-Judge prompt used for comparing model and gold answers in the three DateAugBench tasks.
Human Evaluation Context-based resolution Prompt: Who was the chair of Allgemeiner Deutscher Fahrrad-Club in 17/10/2016? Gold Answer: Ulrich Syberg Model Prediction: As of October 17, 2016, the Federal Chairman was Ulrich Syberg Human Annotator Rating:   A LLM-as-Judge Rating:   A Date arithmetic Prompt: What date is 60 days after 05/01/1225? Gold Answer: March 6, 1225 , June 29, 1225 Model Prediction: July 30, 1225 Human Annotator Rating:   B LLM-as-Judge Rating:   B
Table 8: Human evaluation of LLM-as-judge.